OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7522 — Collection in the 70
Scheduled Maintenance Report for Web Cloud
Completed
There was an incident in the department 70 on
SFR collection.

A ticket was opened, investigations from
their side are in progress.


Update(s):

Date: 2012-10-26 16:47:41 UTC
The access is back.

Date: 2012-10-26 16:43:40 UTC
Services relaunched precisely from 3:30 pm

Date: 2012-10-26 09:25:30 UTC
Une fibre suite a des travaux a ete coupe.
A fiber following to woek tasks has been cut.
Orange will be impacted and it will also onto
place.

The restoration is planned for late afternoon.


Impacted NRA:
BRE70
VTG70
SLP70
PSS70
MEL70
LUR70
Posted Oct 26, 2012 - 09:21 UTC